Find the domain of f(x) = 1 - 4x for which f(a) = 3.
Answer by stanbon(75887)

Find the domain of f(x) = 1 - 4x for which f(a) = 3.
-----
f(a) = 1-4a = 3
---
-4a = 2
a = -1/2
---------------
Cheers,
Stan H.
---------------